OpenStudy (anonymous):

Prove that (b^n)-1 is always divisible by (b-1) for all positive integers b and n.

OpenStudy (anonymous):

b>1 as well.

OpenStudy (anonymous):

1- show that it's true for n=1 ===> L.H.S = b^1 -1 = b-1 ===> it's divisib;e by (b-1).

OpenStudy (anonymous):

2. assume it's true for n= k , i.e b^k -1 is divisible by b-1

OpenStudy (anonymous):

3. show that it's true for n =k+1 ===> b^(k+1) is divisible by b-1, ===> b^(k+1) =b* b^k ===> from step 2. b^k is divisible by b-1 ===> b*b^kk is divisible by b-1 ===> b^(k+1) is divisible by b-1 ===> it's true for n>=1

OpenStudy (anonymous):

this is what is called as proof using mathmatical induction.

OpenStudy (anonymous):

hope that helps.

OpenStudy (anonymous):

Nice proof.

OpenStudy (anonymous):

great then.